### v1

# DeFi Liquidation Risk Modeling Using the Reflection Principle for Zero-Drift Brownian Motion

Abstract
In this paper, we propose an analytical method to compute the collateral liquidation probability in decentralized finance (DeFi) stablecoin single-collateral lending. Our approach models the collateral exchange rate as a zero-drift geometric Brownian motion, converts it into a regular zero-drift Brownian motion, and employs the reflection principle to derive the liquidation probability. Unlike most existing methods that rely on computationally intensive simulations such as Monte Carlo, our formula provides a lightweight, exact solution. This advancement offers a more efficient alternative for risk assessment in DeFi platforms.
1 Introduction
In recent years, decentralized finance (DeFi) lending platforms have gained significant popularity as innovative alternatives to traditional financial services. These platforms enable users to borrow and lend digital assets in a transparent manner. One of the key concerns in DeFi lending is the risk of collateral liquidation, which can lead to substantial financial losses for lenders and borrowers alike. Collateral liquidation occurs when the collateral value falls below the liquidation threshold. Therefore, developing robust methods to assess and quantify the likelihood of liquidation events is essential to ensure the stability and security of DeFi ecosystems.
Predicting liquidation events on DeFi lending platforms is inherently challenging due to the high volatility of cryptocurrency prices.
Furthermore, existing methods for estimating liquidation probabilities often rely on computationally intensive techniques such as Monte Carlo simulations. Although these approaches can provide valuable information, they tend to be resource-intensive and time-consuming, limiting their practicality for real-time risk assessment and decision making. Developing analytical solutions that are accurate and computationally efficient remains an important goal in the field.
Modeling collateral exchange rates as zero-drift Wiener processes offers a significant advantage in terms of computational efficiency and accuracy. This approach allows for the derivation of explicit analytical formulas, eliminating the need for resource-intensive simulations, such as Monte Carlo methods. As a result, the solution becomes not only faster to compute, but also more precise, providing reliable estimates of liquidation probabilities without the errors typically associated with numerical simulations. This makes the method particularly suitable for real-time risk assessment and decision making in the volatile environment of cryptocurrency markets.
2 Deriving Analytical Solution
2.1 Modeling Collateral Exchange Rates as Zero-Drift Brownian Motion
Liquidation occurs when the value of the collateral drops below the liquidation threshold. Collateral value changes due to exchange rates, interest rates and the loan value, although exchange rate volatility diminishes the relevance of interest rates, allowing them to be disregarded in the context of mathematical modeling. We will model loans of stablecoins; therefore, we can assume that the loan value is constant and only model the collateral exchange rates. Also, we will model single-collateral loans.
Both DeFi and CeFi use relative change, not absolute change. This means that exchange rates are multiplied when they change instead of something being added to them or subtracted from them. This means that we can model the exchange rate as the following SDE:
$dS_{t}=\sigma S_{t}dW_{t}$
where $\sigma$ is the standard deviation or volatility, $S_{t}$ is the exchange rate and $W_{t}$ is a Wiener process. This is the equation of zero-drift geometric Brownian motion (GBM) [8] [7]. It is not convenient to work with this equation. So we divide it by $S_{t}$:
${\frac{dS_{t}}{S_{t}}}=\sigma dW_{t}$
Then we integrate both sides on $[0,t]$:
[…]
We get:
$ln\left(S_{t}\right)-ln\left(S_{0}\right)=\sigma W_{t}-\sigma W_{0}=\sigma W_{t}$
We can define $X_{t}$:
$X_{t}=ln\left(S_{t}\right)=\sigma W_{t}+ln\left(S_{0}\right)$
Now it’s a Brownian motion and a simple additive process:
$X_{t+\Delta t}=X_{t}+\sigma\sqrt{\Delta t}Z,\quad Z\sim\mathcal{N}\left(0,1\right)$
Or
$dX_{t}=\sigma dW_{t}$
2.2 Applying the Reflection Principle
Since we disregard interest rates, we can define the liquidation threshold and the exchange rate at which collateral value reaches the liquidation threshold to be constant. Therefore, we can use the reflection principle.
The reflection principle tells us about the probability of a Wiener process crossing some value in some time from now [6]. Here it is:
$\mathbb{P}\left(\underset{0\leq s\leq t}{inf}W_{s}\leq a\right)=2\mathbb{P}\left(W_{t}\leq a\right)$
To ensure applicability to $X$, we can modify it. The starting exchange rate is $S_{0}$, so $X_{0}=ln(S_{0})$. Similarly, $X_{liquidation}=ln(S_{liquidation})$. We can now define $a$:
$a=X_{\text{liquidation}}-X_{0}=ln\left(S_{\text{liquidation}}\right)-ln\left(S_{0}\right)=ln\left(\frac{S_{\text{liquidation}}}{S_{0}}\right)$
After introducing standard deviation $\sigma$ we get:
$\mathbb{P}\left(\text{liquidation}\right)=\mathbb{P}\left(\underset{0\leq s\leq t}{inf}X_{s}\leq a\right)=2\mathbb{P}\left(X_{t}\leq a\right)=2\Phi\left(\frac{a}{\sigma\sqrt{t}}\right)=2\Phi\left(\frac{ln\left(\frac{S_{\text{liquidation}}}{S_{0}}\right)}{\sigma\sqrt{t}}\right)$
Where $\Phi$ is the cumulative distribution function of standard normal distribution. Now we have a complete analytical solution to our problem.
3 Simulating Collateral Exchange Rates
Monte Carlo simulation [3] [4] [5] can be used to validate the theoretical formula derived earlier. To simulate collateral exchange rate paths, we model the evolution of the price as a stochastic process with Gaussian noise using a small timestep $dt$. There are two common ways to discretize the GBM for simulation:
  • Euler–Maruyama approximation: This linear method updates the exchange rate using the first-order approximation of the SDE, assuming
    $dS\sim\sigma S_{t}\sqrt{dt}Z,\quad Z\sim\mathcal{N}\left(0,1\right)$
    which leads to:
    $S_{t+dt}=S_{t}+\sigma S_{t}\sqrt{dt}Z$
  • Exact GBM solution: Alternatively, we can use the closed-form solution to the GBM SDE, which yields multiplicative updates:
    $S_{t+dt}=S_{t}\exp\left(\sigma\sqrt{dt}Z\right)$
For each simulated price path, we check whether the price crosses the liquidation threshold at any time. The simulated liquidation probability is then calculated as:
$\mathbb{P}(\text{liquidation})_{\text{Monte Carlo}}=\frac{N_{\text{liquidated}}}{N}$
where $N_{\text{liquidated}}$ is the number of paths that were liquidated and $N$ is the total number of simulated paths.
4 Simulation Results and Validation
[…]
Figure 1: Probability calculated by the reflection principle and probability calculated by Monte Carlo.
Days Theory GBM Euler-Maruyama Abs. Dif. GBM Abs. Dif. Euler-Maruyama
3 0.003288 0.00184 0.00231 0.001448 0.000978
7 0.054313 0.03556 0.04241 0.018753 0.011903
14 0.173605 0.13239 0.15020 0.041215 0.023405
30 0.352608 0.30196 0.33991 0.050648 0.012698
60 0.510997 0.46529 0.51849 0.045707 0.007493
90 0.591495 0.55254 0.61697 0.038955 0.025475
120 0.642095 0.60637 0.67139 0.035725 0.029295
Table 1: Comparison of liquidation probabilities and their absolute differences from the theoretical values
Figure 1 shows the liquidation probability within the given time in days calculated using the formula derived earlier, the probabilities calculated by Monte Carlo with one hundred thousand simulations for each method and the absolute difference between theory and simulation for each method. The starting exchange rate is 1500 and the exchange rate at which the liquidation happens is 1212. Standard deviation or annualized volatility $\sigma=0.8$. From the results we see that the formula derived closely aligns with the simulations. From the figure we see that the difference doesn’t rise with the increase in time.
5 Applications in DeFi Protocol Design
One of the most direct and impactful applications of our analytical liquidation model lies in its integration with protocols like Aave, one of the leading DeFi lending platforms. In Aave [1] [2], each borrowing position is associated with a health factor, which determines how close it is to liquidation. This health factor is based on the current loan-to-value (LTV) ratio, asset prices, and collateral thresholds. However, it does not explicitly represent the probability that a position will be liquidated within a given time frame. By incorporating our closed-form liquidation probability, Aave could offer an additional dimension to risk assessment: a forward-looking probability metric that captures expected liquidation risk over the next N days, based on historical volatility.
This model is directly useful for developers building Aave-related infrastructure, especially in automation and risk monitoring. Liquidation bots can use the formula to prioritize positions based on short-term liquidation probability, reducing unnecessary checks and API calls. Off-chain dashboards and monitoring tools could show not only health factors but also estimated liquidation risks over time windows. This helps DeFi analysts and risk teams detect emerging threats across collateral pools without relying on simulation-heavy backends. Given its low computational cost, the formula fits well in lightweight scripts and browser-based trackers — from validator dashboards to liquidation queue managers.
6 Conclusion
In this paper, we derived and validated a closed-form analytical formula for computing liquidation probability in DeFi lending protocols. By modeling the evolution of cryptocurrency prices as a zero-drift geometric Brownian motion and applying the reflection principle from probability theory, we were able to express the likelihood of liquidation in a compact, non-simulation-based form. This method avoids the computational overhead of traditional Monte Carlo simulations while retaining strong accuracy across various time horizons. Through a thorough numerical comparison, we showed that the theoretical estimates match closely with empirical results from both Euler–Maruyama and exact GBM-based simulations, with error remaining stable and small.
Our results have direct practical implications. The analytical approach we propose enables real-time liquidation risk analysis, which can be particularly useful for protocol designers, liquidation bots, and institutional risk managers operating in volatile DeFi markets. It also provides a mathematically grounded foundation for designing low-latency monitoring systems that do not depend on high-throughput simulations. While the current model assumes zero drift, constant volatility and constant loan value, future extensions could incorporate stochastic volatility, user leverage, and dynamically adjusting liquidation thresholds. This work demonstrates how well-established mathematical principles can bring analytical clarity and computational efficiency to modern decentralized financial systems, contributing toward more robust and transparent risk assessment frameworks.
Author Contribution
CRediT: Conceptualization: T.B., G.V.; Data curation: T.B., G.V.; Formal analysis: T.B.; Funding acquisition: No funding was received for this work; Investigation: T.B., G.V.; Methodology: T.B.; Project administration: T.B., G.V.; Resources: T.B., G.V.; Software: G.V.; Supervision: T.B., G.V.; Validation: T.B., G.V.; Visualization: T.B, G.V.; Writing - original draft: T.B.; Writing - review and editing: T.B., G.V.